OpenStudy (anonymous):

The directions for mixing an insect spray say to use 1 1/2 ounces of chemical in each gallon of water. How many ounces of chemical should be mixed with 4 3/5 gallons of water? First, estimate the answer to the application problem. The find the exact answer.

OpenStudy (anonymous):

For 1 gallon of water = 1(1/2) ounces of chemicals For 2 gallon of water = 2*1(1/2) ounces of chemicals Thus, for n gallons of water = n* 1(1/2) ounces of chemicals

OpenStudy (anonymous):

Alright so what is the answer

OpenStudy (anonymous):

For 4(3/5) gallons of water = 4(3/5) *1(1/2) ounces of chemicals.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is the answer 69/10

OpenStudy (anonymous):

@sangya21 is this the answer

OpenStudy (anonymous):

correct. what will be the exact answer?

OpenStudy (anonymous):

6 9/10

OpenStudy (anonymous):

69/10 = 6.9